# Definability of mad families of vector spaces and two local Ramsey theories

### v4

@@ 5.1. Block-splitting families
    Consequently, if $f$ and $g$ are incompatible, then so are $\eta_{f}$ and $\eta_{g}$.
@@
By Lemma 2.3, $\{\langle B^{\prime}\rangle\cap\langle D\rangle:B^{\prime}\in\mathcal{B}\}$ is an almost disjoint family of block subspaces of $\langle D\rangle$. By Lemma 5.3, $|\mathcal{B}|<\mathfrak{bs}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$, so there exists some $A_{\alpha}^{\prime}\leq D$ such that $A_{\alpha}^{\prime}$ is almost disjoint from every element of $\mathcal{B}$. Note that $A_{\alpha}^{\prime}$ may be chosen to be a block subspace as every infinite-dimensional subspace contains a block subspace.
@@
    $\langle A_{\alpha}^{\prime}\rangle\setminus Z_{\xi}^{\sigma_{\alpha}(\xi)}$ is small.
  2. (2)
Proof.
  1. (1)
    By Lemma 5.4, $\langle A_{\alpha}\rangle\cap\langle A\rangle$ is very small for each $A\in\mathcal{F}_{\xi}$, so $\langle A_{\alpha}^{\prime}\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}$ is small.
  2. (2)
    Let $\beta+n<\alpha$, where $\beta$ is a limit ordinal. If $A_{\beta+n}\in\mathcal{B}_{0}$, then $A_{\alpha}$ is almost disjoint from $A_{\beta+n}$ by construction, so we are done. Now assume that $A_{\beta+n}\notin W$. Let $\xi$ be the least ordinal such that $\sigma_{\beta}(\xi)\neq\sigma_{\alpha}(\xi)$. By induction hypothesis, we have that $W:=\langle A_{\beta}^{\prime}\rangle\setminus Z_{\xi}^{\sigma_{\beta}(\xi)}=\langle A_{\beta}^{\prime}\rangle\setminus Z_{\xi}^{1-\sigma_{\alpha}(\xi)}$ is small (and hence very small). Since $A_{\beta+n}\leq A_{\beta}^{\prime}$, we have that for some small set $W^{\prime}$:
We complete the induction by letting $\{A_{\alpha+n}:n<\omega\}$ be any infinite disjoint family below $A_{\alpha}^{\prime}$. We note that $\mathcal{A}$ is a mad family – otherwise, if $B$ is almost disjoint from every element of $\mathcal{A}$, then $B\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$ for all $\alpha<\mathfrak{c}$. Let $\alpha$ be any limit ordinal such that $B_{\alpha}=B$, and we have that $A_{\alpha}\leq B$, a contradiction. It remains to show that $\mathcal{A}$ is full.
Let $Y\subseteq E$, and suppose that $Y$ is $\mathcal{H}(\mathcal{A})$-dense below some $B\in\mathcal{H}(\mathcal{A})$. Since $\mathcal{H}(\mathcal{A})=\bigcup_{\alpha<\mathfrak{c}}\mathcal{H}(\mathcal{A}_{\alpha})$, we may let $\alpha$ be large enough so that $(Y_{\alpha},B_{\alpha})=(Y,B)$ and $B\in\mathcal{H}(\mathcal{A}_{\alpha})$. This means that in the induction step, we have either case (b) or (c). Observe that (c) is not possible – $A_{\alpha}^{\prime}\leq B_{\alpha}$ is almost disjoint from every element of $\mathcal{C}_{\alpha}$, and $A_{\alpha}^{\prime}\in\mathcal{H}(\mathcal{A})$ as $A_{\alpha}$ is compatible with $A_{\alpha+n}$ for all $n$. However, there is no $B^{\prime}\leq A_{\alpha}^{\prime}$ in which $\langle B^{\prime}\rangle\subseteq Y_{\alpha}$, contradicting that $Y$ is $\mathcal{H}(\mathcal{A})$-dense below $B$. This means that case (b) must have occurred in this induction step. Since $B_{\alpha^{\prime}}$ is compatible with $A_{\alpha+n}$ for all $n$, $B_{\alpha^{\prime}}\in\mathcal{H}(\mathcal{A})$ and $\langle B_{\alpha}^{\prime}\rangle\subseteq\langle Y_{\alpha}\rangle$, as desired. ∎
5.3. Completely separable mad family of block subspaces
We dedicate this section to proving Theorem 1.6.
Definition 5.6.
A mad family $\mathcal{A}\subseteq E^{[\infty]}$ is completely separable if for all $Y\subseteq E$ such that $\langle Y\rangle\in\mathcal{I}^{++}(\mathcal{A})$, there exists some $A\in\mathcal{A}$ such that $\langle A\rangle\subseteq\langle Y\rangle$.
Proof of Theorem 1.6.
Let $\{Z_{\alpha}:\alpha<\mathfrak{bs}\}$ be a family of subsets of vectors satisfying the conclusions of Lemma 5.5. We write $Z_{\alpha}^{0}:=Z_{\alpha}$ and $Z_{\alpha}^{1}:=Z_{\alpha}^{c}$ for all $\alpha$. With this, given any almost disjoint family $\mathcal{A}\subseteq E^{[\infty]}$ and $B\in\mathcal{H}^{-}(\mathcal{A})$, there exist $\alpha<\mathfrak{bs}$ and $\tau_{B}^{\mathcal{A}}\in 2^{\alpha}$ such that:
  1. (1)
  2. (2)
Observe that for a fixed almost disjoint family $\mathcal{A}$, $\tau_{B}^{\mathcal{A}}\in 2^{<\mathfrak{bs}}$ is unique, and if $B^{\prime}\leq B$ and $B^{\prime}\in\mathcal{H}^{-}(\mathcal{A})$, then $\tau_{B}^{\mathcal{A}}\subseteq\tau_{B^{\prime}}^{\mathcal{A}}$ We may need that for any $Y\subseteq E$, there are cofinally many $\alpha<\mathfrak{c}$ such that $Y_{\alpha}=Y$. We shall define our mad family $\mathcal{A}=\{A_{\alpha}:\alpha<\mathfrak{c}\}$ recursively, where $\mathcal{A}_{\alpha}:=\{A_{\beta}:\beta<\alpha\}$, along with an injective sequence $\{\sigma_{\alpha}:\alpha<\mathfrak{c}\}$ of elements of $2^{<\mathfrak{bs}}$ such that the following holds:
  1. (1)
  2. (2)
  1. (1)
    If $Y_{\alpha}\in\mathcal{I}^{+}(\mathcal{A})$, let $C_{\emptyset}\in\mathcal{H}^{-}(\mathcal{A})$ such that $\langle C_{\emptyset}\rangle\subseteq Y_{\alpha}$. This is possible by Lemma 5.5. If $Y_{\alpha}\notin\mathcal{I}^{+}(\mathcal{A})$, let $C_{\emptyset}\in\mathcal{H}^{-}(\mathcal{A})$ be any bloc subspace.
  2. (2)
  3. (3)
    $C_{s^{\frown}0},C_{s^{\frown}1}\in\mathcal{H}^{-}(\mathcal{A})$ are such that $\langle C_{s^{\frown}0}\rangle\subseteq\langle C_{s}\rangle\cap Z_{\alpha_{s}}^{0}$ and $\langle C_{s^{\frown}1}\rangle\subseteq\langle C_{s}\rangle\cap Z_{\alpha_{s}}^{1}$ (which both exist by properties (ii) and (iii) of Lemma 5.5).
  1. (1)
    If $s\sqsubseteq t$, then $C_{s}\leq C_{t}$ and $\eta_{s}\sqsubseteq\eta_{t}$.
  2. (2)
    If $s$ and $t$ are incompatible, then so are $\eta_{s}$ and $\eta_{t}$.
  3. (3)
    Consequently, if $f$ and $g$ are incompatible, then so are $\eta_{f}$ and $\eta_{g}$.
$\mathcal{B}$ $:=\mathcal{B}_{0}\cup\mathcal{B}_{1}.$
By Lemma 2.3, $\{\langle B^{\prime}\rangle\cap\langle D\rangle:B^{\prime}\in\mathcal{B}\}$ is an almost disjoint family of block subspaces of $\langle D\rangle$. By Lemma 5.3, $|\mathcal{B}|<\mathfrak{bs}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$, so there exists some $A_{\alpha}^{\prime}\leq D$ such that $A_{\alpha}^{\prime}$ is almost disjoint from every element of $\mathcal{B}$. Note that $A_{\alpha}^{\prime}$ may be chosen to be a block subspace as every infinite-dimensional subspace contains a block subspace.
Claim 9.
  1. (1)
    $\langle A_{\alpha}\rangle\setminus Z_{\xi}^{\sigma_{\alpha}(\xi)}$ is small.
@@ 5.3. Completely separable mad family of block subspaces
    Let $\beta<\alpha$. If $A_{\beta}\in\mathcal{B}_{0}$, then $A_{\alpha}$ is almost disjoint from $A_{\beta}$ by construction, so we are done. Now assume that $A_{\beta}\notin W$. Let $\xi$ be the least ordinal such that $\sigma_{\beta}(\xi)\neq\sigma_{\alpha}(\xi)$. By induction hypothesis, we have that $W:=\langle A_{\beta}\rangle\setminus Z_{\xi}^{\sigma_{\beta}(\xi)}=\langle A_{\beta}\rangle\setminus Z_{\xi}^{1-\sigma_{\alpha}(\xi)}$ is small (and hence very small). Thus, we have that for some small set $W^{\prime}$:
@@
This completes the induction. If $\langle Y\rangle\in\mathcal{I}^{+}(\mathcal{A})$, then $\mathcal{I}^{+}(\mathcal{A}_{\alpha})$ for all $\alpha$, and there exists some $\alpha$ such that $Y_{\alpha}=\langle Y\rangle$. This implies that $A_{\alpha}\subseteq\langle Y\rangle$, as desired. Setting $\langle Y\rangle=\langle B\rangle$ for some $B\in E^{[\infty]}$, this implies that $\mathcal{A}$ is mad. ∎
@@ 6. Topological Ramsey spaces and the abstract Mathias forcing
We dedicate this section to an overview of the axioms of topological Ramsey spaces, and the abstract Mathias forcing developed in [DMN15].
@@ 6.1. Topological Ramsey spaces
We recap the four axioms presented by Todorčević in [T10], which are sufficient conditions for a triple $(\mathcal{R},\leq,r)$ to be a topological Ramsey space. Here, $\mathcal{R}$ is a non-empty set, $\leq$ be a quasi-order on $\mathcal{R}$, and $r:\mathcal{R}\times\omega\to\mathcal{AR}$ is a function. We also define a sequence of maps $r_{n}:\mathcal{R}\to\mathcal{AR}$ by $r_{n}(A):=r(A,n)$ for all $A\in\mathcal{R}$. Let $\mathcal{AR}_{n}\subseteq\mathcal{AR}$ be the image of $r_{n}$ (i.e. $a\in\mathcal{AR}_{n}$ iff $a=r_{n}(A)$ for some $A\in\mathcal{R}$).
@@
    For each $a\in\mathcal{AR}$, let $\operatorname{lh}(a)$ denote the unique $n$ in which $a\in\mathcal{AR}_{n}$. By Axiom A1(3), this $n$ is well-defined.
@@
Note that by axiom A1, we may identify each element $A\in\mathcal{R}$ as a sequence of elements of $\mathcal{AR}$, via the map $A\mapsto(r_{n}(A))_{n<\omega}$. Therefore, we may identify $\mathcal{R}$ as a subset of $\mathcal{AR}^{\mathbb{N}}$.
@@
A triple $(\mathcal{R},\leq,r)$ is said to be a closed triple if $\mathcal{R}$ is a metrically closed subset of $\mathcal{AR}^{\mathbb{N}}$.
A closed triple satisfying $(\mathcal{R},\leq,r)$ satisfying A1-A4 is a topological Ramsey space, as it satisfies the abstract Ellentuck theorem – that is, a subset $\mathcal{X}\subseteq\mathcal{R}$ is Ramsey iff it has the property of Baire relative to the Ellentuck topology. This is the abstract Ellentuck theorem by Todorčević – see Theorem 5.4 of [T10].
We list some commonly used terminologies and notations for topological Ramsey spaces.
  • If $A\leq B$, we say that $A$ is a reduction of $B$.
  • $\mathcal{AR}\mathord{\upharpoonright}A:=\{b\in\mathcal{AR}:\exists n[b\leq_{\mathrm{fin}}r_{n}(A)]\}$.
  • $\mathcal{AR}\mathord{\upharpoonright}[a,A]:=\{b\in\mathcal{AR}:a\sqsubseteq b\wedge b\in\mathcal{AR}\mathord{\upharpoonright}A\}$.
Mathias’ original proof of his celebrated theorem (i.e. Theorem 1.7 for $\omega]^{\omega}$) uses the (original) Mathias forcing localised to a coideal $\mathcal{H}$. As we shall use a similar approach to prove Theorem 1.7, it is necessary to extend the notion of a coideal to topological Ramsey spaces.
Definition 6.2 (Definition 3.1, [DMN15]).
@@
A notion of semiselectivity was also given as Definition 3.4-3.7 in [DMN15]. We propose a slight variant of this definition.
@@
Let $(\mathcal{R},\leq,r)$ be an A2-space, and let $\mathcal{H}\subseteq\mathcal{R}$ be a semicoideal. Let $A\in\mathcal{H}$ and $a\in\mathcal{AR}\mathord{\upharpoonright}A$.
@@
    A family of subsets $\vec{\mathcal{D}}=\{\mathcal{D}_{b}\}_{b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]}$ is dense open below $[a,A]$ in $\mathcal{H}$ if for all $b\in\mathcal{AR}\mathord{\upharpoonright}[b,A]$, $\mathcal{D}_{b}$ is a $\leq$-downward closed subset of $\mathcal{H}\mathord{\upharpoonright}[b,A]$, and for all $B\in\mathcal{H}\mathord{\upharpoonright}[b,A]$, there exists some $C\in\mathcal{H}\mathord{\upharpoonright}[b,B]$ such that $C\in\mathcal{D}_{b}$.
@@
This is stated as Lemma 3.11 of [Y24]. We replicate the proof here.
@@
Note that $\mathcal{R}$ is a coideal by definition. Fix some $A\in\mathcal{R}$ and $a\in\mathcal{AR}\mathord{\upharpoonright}A$. Suppose that $\vec{\mathcal{D}}=\{\mathcal{D}_{b}\}_{b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]}$ is dense open below $[a,A]$. We shall define a fusion sequence $(A_{n})_{n<\omega}$ in $[a,A]$, with $a_{n+1}:=r_{\operatorname{lh}(a)+n+1}(A_{n})$, such that $A_{n+1}\in[a_{n+1},A_{n}]$: Let $A_{0}:=A$, and suppose that $A_{n}$ has been defined. Let $\{b_{i}:i<N\}$ enumerate the set of all $b\in\mathcal{AR}\mathord{\upharpoonright}A_{n}$ such that $a\sqsubseteq b$ and $b\leq_{\mathrm{fin}}a_{n+1}$. Let $A_{n+1}^{0}:=A_{n}$. If $A_{n+1}^{i}\in[a_{n+1},A_{n}]$ has been defined, let $B_{n+1}\in\mathcal{D}_{b_{i}}$ be such that $B_{n+1}\in[b_{i},A_{n+1}^{i}]$, which exists as $\mathcal{D}_{b_{i}}$ is dense open in $b_{i},A]$. By A3, we then let $A_{n+1}^{i+1}\in[a_{n+1},A_{n}^{i}]$ such that $b_{i},A_{n+1}^{i+1}]\subseteq[b_{i},B_{n+1}]$. We complete the induction by letting $A_{n+1}:=A_{n+1}^{N}$. Let $B$ be the limit of the fusion sequence $(A_{n})_{n<\omega}$, and we have that $B$ diagonalises $\vec{\mathcal{D}}$. ∎

### v5

@@ 5.3. Completely separable mad family of subspaces
    Consequently, $f\neq g$, then $\eta_{f}$ and $\eta_{g}$ are incompatible.
@@
One may show, in the same way of proving Claim 8, that $|\mathcal{B}|<\mathfrak{bs}$. By Lemma 2.3, the set:
$\{\langle B^{\prime}\rangle\cap\langle D\rangle:B^{\prime}\in\mathcal{B}\text{ and }\langle B^{\prime}\rangle\cap\langle D\rangle\text{ is infinite-dimensional}\}$
is an almost disjoint family of block subspaces of $\langle D\rangle$. Since $D\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$, we have that $\langle D\rangle$ is not in this set. By Lemma 5.3, $|\mathcal{B}|<\mathfrak{bs}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$, so there exists some $A_{\alpha}\leq D$ such that $A_{\alpha}$ is almost disjoint from every element of $\mathcal{B}$.
@@
    For all $\xi\in\operatorname{dom}(\sigma_{\alpha})$, $\langle A_{\alpha}\rangle\setminus Z_{\xi}^{\sigma_{\alpha}(\xi)}$ is small.
@@
    Let $\beta<\alpha$. If $A_{\beta}\in\mathcal{B}_{0}$, then $A_{\alpha}$ is almost disjoint from $A_{\beta}$ by construction, so we are done. Now assume that $A_{\beta}\notin\mathcal{B}_{0}$. Let $\xi$ be the least ordinal such that $\sigma_{\beta}(\xi)\neq\sigma_{\alpha}(\xi)$. By induction hypothesis, we have that $W:=\langle A_{\beta}\rangle\setminus Z_{\xi}^{\sigma_{\beta}(\xi)}=\langle A_{\beta}\rangle\setminus Z_{\xi}^{1-\sigma_{\alpha}(\xi)}$ is small (and hence very small). Thus, we have that for some small set $W^{\prime}$:
@@
This completes the induction. If $U\in\mathcal{I}^{+}(\mathcal{A})$, then $\mathcal{I}^{+}(\mathcal{A}_{\alpha})$ for all $\alpha$, and there exists some $\alpha$ such that $U_{\alpha}=U$. This implies that $\langle A_{\alpha}\rangle\subseteq U$, as desired. ∎
@@ 6. Topological Ramsey spaces and the abstract Mathias forcing
We dedicate this section to an overview of the axioms of topological Ramsey spaces, and the abstract Mathias forcing developed in [6].
@@ 6.1. Topological Ramsey spaces
We recap the four axioms presented by Todorčević in [27], which are sufficient conditions for a triple $(\mathcal{R},\leq,r)$ to be a topological Ramsey space. Here, $\mathcal{R}$ is a non-empty set, $\leq$ a quasi-order on $\mathcal{R}$, and $r:\mathcal{R}\times\omega\to\mathcal{AR}$ a surjective function. We also define a sequence of maps $r_{n}:\mathcal{R}\to\mathcal{AR}$ by $r_{n}(A):=r(A,n)$ for all $A\in\mathcal{R}$. Let $\mathcal{AR}_{n}\subseteq\mathcal{AR}$ be the image of $r_{n}$ (i.e. $a\in\mathcal{AR}_{n}$ iff $a=r_{n}(A)$ for some $A\in\mathcal{R}$).
@@
    For each $a\in\mathcal{AR}$, let $\operatorname{lh}(a)$ denote the unique $n$ in which $a\in\mathcal{AR}_{n}$. By Axiom A1(3), this $n$ is well-defined. Given $a,b\in\mathcal{AR}$, we write $a\sqsubseteq b$ if there exists some $A\in\mathcal{R}$ such that $a=r_{n}(A)$ and $b=r_{m}(A)$ for some $n\leq m$.
@@
Note that by axiom A1, we may identify each element $A\in\mathcal{R}$ as a sequence of elements of $\mathcal{AR}$, via the map $A\mapsto(r_{n}(A))_{n<\omega}$. Therefore, we may identify $\mathcal{R}$ as a subset of $\mathcal{AR}^{\mathbb{N}}$. $\mathcal{AR}^{\mathbb{N}}$ may be equipped with a natural metric topology by considering the first difference metric — that is, if $(a_{n})_{n<\omega},(b_{n})_{n<\omega}\in\mathcal{AR}^{\mathbb{N}}$, then:
$d((a_{n})_{n<\omega},(b_{n})_{n<\omega})=2^{-\min\{n:a_{n}\neq b_{n}\}}.$
@@
A triple $(\mathcal{R},\leq,r)$ is said to be a closed triple if $\mathcal{R}$ is a metrically closed subset of $\mathcal{AR}^{\mathbb{N}}$. A closed triple $(\mathcal{R},\leq,r)$ is an A2-space if $(\mathcal{R},\leq,r)$ satisfies the axioms A1-A3.
An A2-space $(\mathcal{R},\leq,r)$ satisfying A4 is a topological Ramsey space, as it satisfies the abstract Ellentuck theorem — that is, a subset $\mathcal{X}\subseteq\mathcal{R}$ is Ramsey iff it has the property of Baire relative to the Ellentuck topology. This is the abstract Ellentuck theorem by Todorčević — see Theorem 5.4 of [27].
Mathias’ original proof of his celebrated theorem in [13] (i.e. Theorem 1.7 for $\omega]^{\omega}$) uses the (original) Mathias forcing localised to a coideal $\mathcal{H}$. As we shall use a similar approach to prove Theorem 1.7, it is necessary to extend the notion of a coideal to topological Ramsey spaces.
Definition 6.2 (Definition 3.1, [6]).
@@
A notion of semiselectivity for a coideal $\mathcal{H}$ was also given as Definition 3.4-3.7 in [6], which says for all $A\in\mathcal{H}$ and a family of dense open sets, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}A$ that diagonalises the family. We shall propose a slightly stronger version of semiselectivity, which instead asks that for all $A\in\mathcal{H}$, $a\in\mathcal{AR}\mathord{\upharpoonright}A$ and family of dense open sets, there exists some $B\in\mathcal{H}\mathord{\upharpoonright}[a,A]$ such that $B$ diagonalises the family.
@@
Let $(\mathcal{R},\leq,r)$ be a closed triple satisfying A1-A4, and let $\mathcal{H}\subseteq\mathcal{R}$ be a coideal. Let $A\in\mathcal{H}$ and $a\in\mathcal{AR}\mathord{\upharpoonright}A$.
@@
    A family of subsets $\vec{\mathcal{D}}=\{\mathcal{D}_{b}\}_{b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]}$ is dense open below $[a,A]$ in $\mathcal{H}$ if for all $b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]$, $\mathcal{D}_{b}$ is a $\leq$-downward closed subset of $\mathcal{H}\mathord{\upharpoonright}[b,A]$, and for all $B\in\mathcal{H}\mathord{\upharpoonright}[b,A]$, there exists some $C\in\mathcal{H}\mathord{\upharpoonright}[b,B]$ such that $C\in\mathcal{D}_{b}$.
@@
This is stated as Lemma 3.11 of [31]. We replicate the proof here.
@@
Note that $\mathcal{R}$ is a coideal by definition. Fix some $A\in\mathcal{R}$ and $a\in\mathcal{AR}\mathord{\upharpoonright}A$. Suppose that $\vec{\mathcal{D}}=\{\mathcal{D}_{b}\}_{b\in\mathcal{AR}\mathord{\upharpoonright}[a,A]}$ is dense open below $[a,A]$. We shall define a fusion sequence $(A_{n})_{n<\omega}$ in $[a,A]$, with $a_{n+1}:=r_{\operatorname{lh}(a)+n+1}(A_{n})$, such that $A_{n+1}\in[a_{n+1},A_{n}]$: Let $A_{0}:=A$, and suppose that $A_{n}$ has been defined. Let $\{b_{i}:i<N\}$ enumerate the set of all $b\in\mathcal{AR}\mathord{\upharpoonright}A_{n}$ such that $a\sqsubseteq b$ and $b\leq_{\mathrm{fin}}a_{n+1}$. Let $A_{n+1}^{0}:=A_{n}$. If $A_{n+1}^{i}\in[a_{n+1},A_{n}]$ has been defined, let $B_{n+1}\in\mathcal{D}_{b_{i}}$ be such that $B_{n+1}\in[b_{i},A_{n+1}^{i}]$, which exists as $\mathcal{D}_{b_{i}}$ is dense open in $b_{i},A]$. By A3, we then let $A_{n+1}^{i+1}\in[a_{n+1},A_{n}^{i}]$ such that $b_{i},A_{n+1}^{i+1}]\subseteq[b_{i},B_{n+1}]$. We complete the induction by letting $A_{n+1}:=A_{n+1}^{N}$. Let $B$ be the limit of the fusion sequence $(A_{n})_{n<\omega}$. Since $\mathcal{R}$ is metrically closed in $\mathcal{AR}^{\mathbb{N}}$, $B\in\mathcal{R}$ is well-defined, and we have that $B$ diagonalises $\vec{\mathcal{D}}$. ∎